# The reliable quantum master equation of the Unruh-DeWitt detector

### v2

@@ V Conclusion
We present a method to estimate the validity range of the quantum Markovian master equation (QMME) for the UDW detector in a general context, particularly without requiring any exact solution for the detector’s evolution. Our key results include reliability conditions (32) for the Markov approximation and applicability conditions (48) for the RWA. Both validity conditions are $g$-dependent, indicating that they hold for $g\rightarrow 0$, consistent with Davies’ theorems. Nevertheless, the specific forms of these conditions depend on the details of the setting of an open system. For the UDW detector model, these could be the spacetime background, the trajectory of the detector, and the type of quantum field being studied. We illustrate this by re-examining the open dynamics of an accelerating UDW detector undergoing the Unruh effect, where the valid conditions narrow the parameter space to ensure the reliability of the prediction of the QMME (23). In particular, we found that the errors caused by the Markov and RWA approximations are bounded in most parameter regions without requiring extremely weak coupling. Only in regions with small detector energy spacing or spatial size do the error terms need extremely weak coupling to suppress them and restore the reliability of the QMME.
@@
Our analysis did not account for potential errors stemming from the Born approximation. This assumption could prove inadequate, particularly when the ”bath” is insufficiently large to disregard the detector’s backreaction (such as when the detector engages with small, non-ideal thermal reservoirs). It would be important to extend our method in future to assess the specific nature of the errors caused by the Born approximation, completing our investigation into the reliability of the QMME.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12475061 and 12075178), the Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSY006), and the Innovation Program for Quantum Science and Technology (2021ZD0302400).
@@ 2. The operator norm:
For convenience, we denote the second term of the error part by $\Delta$. If we expand the commutator in $\Delta$ and consider its Hermitian conjugate, $\Delta$ consists of four parts, which obey exactly the same bound since they differ from each other only in the operator order, which is irrelevant for taking the norm. Therefore, using the triangle inequality and submultiplicativity of the operator norm $\|\cdot\|_{\infty}$ sec2-1 , we have:
@@
Next, we want to calculate the order of $\|\Delta\|_{\infty}$ under the late-time limit $g^{2}t\sim\mathcal{O}(1)$, which in our analysis replaces the more restrictive van Hove limit. To this aim, we first transform the real-time $t$ to a rescaled time $\mathcal{T}=g^{2}t$, so that $\mathcal{T}$ is $\mathcal{O}(1)$, and the inequality (78) becomes:
@@
Further calculations require the specific form of $\mathcal{W}(\tau)$, which, for now, we can assume has a power-law dependence over time:
@@
which gives the bound of $\|\Delta\|_{\infty}$ as:

### v3

@@ V Conclusion
We present a method to estimate the validity range of the QMME for the UDW detector in a general context, particularly without requiring any exact solution for the detector’s evolution. Our key results include reliability conditions (32) for the Markov approximation and applicability conditions (48) for the RWA. Both validity conditions are $g$-dependent, indicating that they hold for $g\rightarrow 0$, consistent with Davies’ theorems. Nevertheless, the specific forms of these conditions depend on the details of the setting of an open system. For the UDW detector model, these could be the spacetime background, the trajectory of the detector, and the type of quantum field being studied. We illustrate this by reexamining the open dynamics of an accelerating UDW detector undergoing the Unruh effect, where the valid conditions narrow the parameter space to ensure the reliability of the prediction of the QMME (23). In particular, we found that the errors caused by the Markov and RWA approximations are bounded in most parameter regions without requiring extremely weak coupling. Only in regions with small detector energy spacing or spatial size do the error terms need extremely weak coupling to suppress them and restore the reliability of the QMME.
@@
Our analysis did not account for potential errors stemming from the Born approximation. This assumption could prove inadequate, particularly when the “bath” is insufficiently large to disregard the detector’s backreaction (such as when the detector engages with small, nonideal thermal reservoirs). It would be important to extend our method in future to assess the specific nature of the errors caused by the Born approximation, completing our investigation into the reliability of the QMME.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 12475061 and No. 12075178), the Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSY006), and the Innovation Program for Quantum Science and Technology (2021ZD0302400).
@@ 2. The operator norm:
For convenience, we denote the second term of the error part by $\Delta$. If we expand the commutator in $\Delta$ and consider its Hermitian conjugate, $\Delta$ consists of four parts, which obey exactly the same bound since they differ from each other only in the operator order, which is irrelevant for taking the norm. Therefore, using the triangle inequality and submultiplicativity of the operator norm $\|\cdot\|_{\infty}$ sec2-1 , we have
@@
Next, we want to calculate the order of $\|\Delta\|_{\infty}$ under the late-time limit $g^{2}t\sim\mathcal{O}(1)$, which in our analysis replaces the more restrictive van Hove limit. To this aim, we first transform the real-time $t$ to a rescaled time $\mathcal{T}=g^{2}t$, so that $\mathcal{T}$ is $\mathcal{O}(1)$, and the inequality (78) becomes
@@
Further calculations require the specific form of $\mathcal{W}(\tau)$, which, for now, we can assume has a power-law dependence over time,
@@
which gives the bound of $\|\Delta\|_{\infty}$ as